OpenStudy (anonymous):

Find the digit in the units place of the integer 3^444 Can someone help me solve this.Its not for a test,Im just trying to gain some extra credit points

OpenStudy (anonymous):

look at the pattern of three to powers

OpenStudy (anonymous):

\[3^0=1\\ 3^1=3\\ 3^2=9\\ 3^3=27\\ 3^4=81\] now look only at the ones place

OpenStudy (anonymous):

\[3^0\to 1\\ 3^1\to 3\\ 3^2\to 9\\ 3^3\to 7\\ 3^4\to 1\] once you see a 1 the pattern will repeat

OpenStudy (anonymous):

ok i get it but how is 3^4=1??

OpenStudy (radar):

That was what the "ones" or units value, 3^4 was 81, 1 is in the unit position 8 was in the tens position.

OpenStudy (anonymous):

@radar Im still lost

OpenStudy (radar):

He was suggesting you study the pattern, noting when the exponent ended in 4 the result was with a 1 in the for the units digit\[3^{4}=81\] \[3^{8}=6561\] \[3^{12}=531441\]\[3^{16}=43046721\] Do you get the pattern? Note that the value of the units digit is a 1........so just maybe we could just say that \[3^{444}=\]a very large value whose units digit is a 1.

OpenStudy (radar):

*"ending in 4 or a multiple of 4"